40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Entry  Wed Jul 7 13:05:07 2021, Paco, DailyProgress, 1418 nm AUX ECDL, New aspheric flexures 
    Reply  Fri Jul 9 11:40:42 2021, Paco, DailyProgress, 1418 nm AUX ECDL, New aspheric flexures IMG_1773.jpeg
       Reply  Wed Jul 14 12:21:08 2021, Paco, DailyProgress, 1418 nm AUX ECDL, ECDL lases... and MZ locked mid-fringe selfhomodyne_1418.jpg
          Reply  Wed Aug 4 11:36:30 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL with 2um AOM tests rf_setup.jpgdiffraction_levels.png
             Reply  Tue Aug 10 09:51:44 2021, Paco, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz rf_setup.jpg
                Reply  Thu Aug 12 11:49:59 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz zeroth_order_contour.pdffirst_order_contour.pdf
                   Reply  Tue Aug 17 11:09:29 2021, rana, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz 
                   Reply  Wed Sep 1 13:12:02 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz 
Message ID: 1918     Entry time: Tue Aug 10 09:51:44 2021     In reply to: 1917     Reply to this: 1920
Author: Paco 
Type: DailyProgress 
Category: 1418 nm AUX ECDL 
Subject: 1419 nm ECDL AOM diffraction at 95 MHz 

[radhika, paco]

We changed the setup to use a low power amplifier rather than the 5W amp from last time. The updated schematic is in Attachment 2. This is in part because 5W is an overkill to drive a fiber AOM which is known to saturate at 0.6 mW of RF input, but also because working with lower power active elements is easier and considerably safer. We dropped the 5W amp. in Rana's office last Friday, and got a ZHL-3A-sma. This little guy gives a max power output of 29.5 dBm (~890 mW) which should be more than enough while using the Marconi as our source (max output +13 dBm).

We hooked the amplifier to the load (AOM) without any couplers or attenuators in between, powered it with +24 VDC and quickly repeated a scan of the source power level while to see any sign of diffraction in the PDs. The result is in Attachment 2. We were a little bit disappointed that there appeared to be no diffraction, so next we tried scanning the RF frequency (it was nominally at 80 MHz) around and we finally succeeded in seeing some diffraction at 95 MHz! Paco thinks the internal fiber coupling made for the design wavelength (2004 nm) is suboptimal at 80 MHz and ~1.4 um wavelength. Therefore, to couple the 1st order back into the fiber, we need to shift the RF frequency to restore the diffraction angle at the cost of potentially not driving the optimal efficiency. An interesting observation made at the same time we saw 1st order light was that the power seemed to drift very slowly (-1%/min), which may have to do with some thermal drift inside the crystal... Our plan is to make a complete characterization of the diffraction efficiency at 1.4 um, and also investigate the slow intensity drifts as a function of RF input. The goal is not so much to understand and fix this last one, but to be able to operate the setup at a point where things are stable for a low frequency, frequency noise measurement.

Attachment 1: rf_setup.jpg  23 kB  Uploaded Tue Aug 10 11:57:09 2021  | Show | Show all
ELOG V3.1.3-